# Reinforcement Learning in Vision: A Survey

### v1

@@ 1 Introduction
In recent years, inspired by these remarkable achievements from LLM, there has been an explosive interest in extending the RL methodologies that proved successful for LLMs to multimodal large models, including Vision-Language Models (VLM) (Zhou et al., 2025a; Zheng et al., 2025; Zhang et al., 2025f), Vision-Language-Action models (VLA) (Lu et al., 2025b; Tan et al., 2025b; Luo et al., 2025; Yuan et al., 2025b), diffusion-based visual generation models (Fan et al., 2023b; Black et al., 2023; Zhou et al., 2025c), and unified multimodal frameworks (Mao et al., 2025; Wang et al., 2024b; 2025a), as shown in Figure 1. Multimodal models such as Gemini 2.5 (DeepMind, 2025) have leveraged RL to align visual-textual reasoning processes and produce outputs with higher semantic coherence and alignment with human judgments. Concurrently, VLA models integrating vision and language with action-oriented outputs have adopted RL to optimize complex sequential decision-making processes in interactive environments, significantly improving task-specific performance in GUI automation (Yuan et al., 2025b; Shi et al., 2025), robotic manipulation (Lu et al., 2025b), and embodied navigation (Kim et al., 2025). The rapid advancement of diffusion-based generative models has further spurred this RL-driven innovation wave. Works like ImageReward (Xu et al., 2023) have introduced reinforcement learning to enhance the semantic alignment and visual quality of generative outputs, refining diffusion-based generation through iterative feedback mechanisms derived from human preferences or automated reward critics. Moreover, unified models that blend multiple tasks, i.e., understanding, and generation, into single architectures (Mao et al., 2025; Jiang et al., 2025b) have increasingly relied on RL-driven fine-tuning, achieving generalization and task transfer previously considered challenging. Despite the substantial progress in integrating reinforcement learning with multimodal large language models, several core challenges remain. These include stabilizing policy optimization under complex reward signals, managing high-dimensional and diverse visual inputs, and designing scalable reward functions that support long-horizon decision-making. Addressing these challenges necessitates methodological innovations in both algorithmic design and evaluation protocols.
@@ 3 Reinforcement Learning in Vision
Frameworks(§3.1.1), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  RePIC (Oh et al., 2025), GoalLadder (Zakharov & Whiteson, 2025), Drive-R1 (Li et al., 2025b),
VLM-R1 (Shen et al., 2025), GRPO-CARE (Chen et al., 2025b), Q-Ponder (Cai et al., 2025c),
@@
ProxyThinker (Xiao et al., 2025b), Jigsaw-R1 (Wang et al., 2025i), SRPO (Wan et al., 2025),
R1-Onevision (Yang et al., 2025e), Reason-RFT (Tan et al., 2025a), etc. , cause_leaf, text width=38em ] ] [ Spatial & 3D Perception
@@
Bindgpt (Zholus et al., 2025), Scene-R1 (Yuan et al., 2025c), Perception-R1 (Xiao et al., 2025a),
VisRL (Chen et al., 2025h), ViCrit (Wang et al., 2025f), etc. , cause_leaf, text width=38em ] ] [ Image Reasoning (§3.1.3), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Think about Image, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g.,  SVQA-R1 (Wang & Ling, 2025), VL-GenRM (Zhang et al., 2025c), RACRO (Gou et al., 2025)
, EasyARC (Unsal & Akkus, 2025), STAR-R1 (Li et al., 2025d), Visionary-R1 (Xia et al., 2025),
UniVG-R1 (Bai et al., 2025), EchoInk-R1 (Xing et al., 2025), WeThink (Yang et al., 2025a),
@@
Mm-eureka (Meng et al., 2025), etc. , cause_leaf, text width=38em ] ] [ Think with Image, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g.,  Visual Planning (Xu et al., 2025b), GRIT (Fan et al., 2025), VILASR (Wu et al., 2025a),
@@
DeepEyes (Zheng et al., 2025), TACO (Kan et al., 2025), VRAG-RL (Wang et al., 2025e),
VisTA (Huang et al., 2025),TGI (Chern et al., 2025), Chain-of-Focus (Zhang et al., 2025f),
Openthinkimg (Su et al., 2025b), etc. , cause_leaf, text width=38em ] ] ] [ Video Reasoning(§3.1.4), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  VQ-Insight (Zhang et al., 2025g), TW-GRPO (Dang et al., 2025), EgoVLM (Vinod et al., 2025),
VAU-R1 (Zhu et al., 2025b), DeepVideo-R (Park et al., 2025a), TimeMaster (Zhang et al., 2025d),
VideoR1 (Feng et al., 2025), etc. , cause_leaf, text width=38em ] ] ] [ Visual Generation(§3.2), color=lightplum!100, fill=lightplum!100, text=black [ Image Generation(§3.2.1), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  ImageReward (Xu et al., 2023), ReasonGen-R1 (Zhang et al., 2025h), FocusDiff (Pan et al., 2025),
@@
Black et al., (Black et al., 2023), DiffusionDPO (Wallace et al., 2024),
Diffusion-KTO (Li et al., 2024b), DiffusionRPO (Gu et al., 2024), Miao et al., (Miao et al., 2024b),
Parrot (Lee et al., 2024), DPG-T2I (Wei et al., 2024), RPO (Miao et al., 2024a),
RePrompt (Wu et al., 2025b), RLCM (Oertell et al., 2024), B2-DiffuRL (Hu et al., 2025a),
DiffPPO (Xiao et al., 2024), Simplear(Wang et al., 2025c), LLaVA-Reward (Zhou et al., 2025b),etc. , detect_leaf, text width=38em ] ] [ Video Generation(§3.2.2), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  DanceGRPO (Xue et al., 2025), InfLVG (Fang et al., 2025), Phys-AR (Lin et al., 2025),
VideoReward (Liu et al., 2025a), TeViR (Chen et al., 2025d), GAPO (Zhu et al., 2025a),
InfLVG (Fang et al., 2025), Instructvideo (Yuan et al., 2024), Videoscore (He et al., 2024b),
@@
Nabla-R2D3 (Liu et al., 2025c), etc. , detect_leaf, text width=38em ] ] ] [ Unified Model(§3.3), color=harvestgold!100, fill=harvestgold!100, text=black [ Unified RL(§3.3.1), color=harvestgold!100, fill=harvestgold!60, text=black [ e.g.,  UniRL (Mao et al., 2025), CoRL (Jiang et al., 2025b), SelfTok (Wang et al., 2025a),
Hermesflow (Yang et al., 2025b), etc. , detect_harvestgold, text width=38em ] ] [ Task-specific RL(§3.3.2), color=harvestgold!100, fill=harvestgold!60, text=black [ e.g.,  Vargpt-v1. 1 (Zhuang et al., 2025), Emu3 (Wang et al., 2024b), X-Omni (Geng et al., 2025), etc. , detect_harvestgold, text width=38em ] ] ] [ Vision Language Action
@@
UIShift (Gao et al., 2025b), AgentCPM-GUI (Zhang et al., 2025i), MobileGUI-RL (Shi et al., 2025),
ProgRM (Zhang et al., 2025a), Mobile-R1 (Gu et al., 2025), GTA1 (Yang et al., 2025d),
@@
VLN-R1 (Qi et al., 2025), Flare (Hu et al., 2024a), etc. , detect_burlywood, text width=38em ] ] [ Visual Manipulation(§3.4.3), color=burlywood!100, fill=burlywood!60, text=black [ e.g.,  TGRPO (Chen et al., 2025g), RFTF (Shu et al., 2025), RLVLA (Liu et al., 2025b),
@@
RIPT-VLA (Tan et al., 2025b), ReinBot (Zhang et al., 2025b),
@@ 3.1.1 Conventional RL-based MLLMs
We refer to conventional RL-based MLLMs as approaches that apply reinforcement learning primarily to align a vision–language backbone with verifiable, task-level rewards, without explicitly modeling multi-step chain-of-thought reasoning. Typical works RePIC (Oh et al., 2025), GoalLadder (Zakharov & Whiteson, 2025), Drive-R1 (Li et al., 2025b) and VLM-R1 (Shen et al., 2025) replace preference models with deterministic validators (e.g., exact-match, IoU, BLEU) and optimize the policy by GRPO/PPO variants under a KL regulariser. This design yields stable value-free training, improves zero-shot robustness on captioning, grounding and autonomous-driving benchmarks, and substantially reduces the annotation cost typically incurred by supervised fine-tuning.
@@ 3.1.3 Image Reasoning
Think about Image. Early think about image works for spatial VQA employs view-consistent or transformation-invariant objectives, such as SVQA-R1 (Wang & Ling, 2025) and STAR-R1 (Li et al., 2025d). VL-GenRM (Zhang et al., 2025c) and RACRO (Gou et al., 2025) refine preference data or caption rewards to curb hallucinations. Benchmark-oriented efforts such as EasyARC (Unsal & Akkus, 2025) offer procedurally generated, fully verifiable tasks that suit outcome-based reinforcement learning. To mitigate shortcut reliance and improve generality, Visionary-R1 (Xia et al., 2025) enforces image interpretation before reasoning, whereas UniVG-R1 (Bai et al., 2025) unifies referring, captioning, and detection by coupling a grounding corpus with GRPO fine-tuning. Extensions such as EchoInk-R1 (Xing et al., 2025) further enrich visual reasoning by integrating audio–visual synchrony under GRPO optimization. Meanwhile, curriculum-driven frameworks—WeThink, G1, GThinker, and Observe-R1 progressively increase task complexity or introduce re-thinking cues (e.g., difficulty ladders, multimodal format constraints) to cultivate deeper and more structured reasoning capabilities in MLLMs. These methods show that language only RL with well-designed visual correctness rewards significantly improves model accuracy, robustness, and out-of-distribution performance.
@@ 3.1.4 Video Reasoning
Video reasoning extends the capabilities of MLLMs to process temporal dynamics, requiring not only spatial perception but also sequential understanding and causal inference. Recent works in this domain have proposed diverse approaches to tackle complex reasoning over video inputs. For instance, VQ-Insight (Zhang et al., 2025g) introduces a hierarchical reward design and self-consistency voting mechanism tailored to the question–answering process over long videos. TW-GRPO (Dang et al., 2025) combines token wise credit assignment with GRPO-style optimization to improve fine-grained temporal alignment between textual reasoning and video evidence. Meanwhile, several R1-style frameworks have been developed to unlock video understanding in complex real-world or egocentric settings. EgoVLM (Vinod et al., 2025) and VAU-R1 (Zhu et al., 2025b) focus on egocentric video reasoning with visual memory and utility-based rewards. DeepVideo-R (Park et al., 2025a) integrates dense video encodings and external reward functions to supervise long-horizon reasoning. TimeMaster (Zhang et al., 2025d) explicitly structures temporal abstraction and reasoning via curriculum learning, while VideoR1 (Feng et al., 2025) proposes a scalable RL framework for video-based QA tasks across multiple domains. Collectively, these works highlight the importance of aligning temporal representations with language trajectories through reinforcement learning, paving the way for robust and generalizable video reasoning agents.
@@ 3.2 Visual Generation
Figure 5: Three reward paradigms for RL-based image generation. (a) Human-Centric Preference Optimization: aligns outputs with human aesthetic scores (HPS (Wu et al., 2023b), ImageReward (Xu et al., 2023)); (b) Multimodal Reasoning-Based Evaluation: scores images via multimodal reasoning consistency (UnifiedReward (Wang et al., 2025h), PARM (Guo et al., 2025d)); (c) Metric-Driven Objective Optimization: minimizes task-specific quantitative metrics such as FID and IoU.

### v2

@@ 1 Introduction
In recent years, inspired by these remarkable achievements from LLM, there has been an explosive interest in extending the RL methodologies that proved successful for LLMs to multimodal large models, including Vision-Language Models (VLM) (Zhou et al., 2025a; Zheng et al., 2025; Zhang et al., 2025g), Vision-Language-Action models (VLA) (Lu et al., 2025b; Tan et al., 2025b; Luo et al., 2025; Yuan et al., 2025b), diffusion-based visual generation models (Fan et al., 2023b; Black et al., 2023; Zhou et al., 2025c), and unified multimodal frameworks (Mao et al., 2025; Wang et al., 2024b; 2025a), as shown in Figure 1. Multimodal models such as Gemini 2.5 (DeepMind, 2025) have leveraged RL to align visual-textual reasoning processes and produce outputs with higher semantic coherence and alignment with human judgments. Concurrently, VLA models integrating vision and language with action-oriented outputs have adopted RL to optimize complex sequential decision-making processes in interactive environments, significantly improving task-specific performance in GUI automation (Yuan et al., 2025b; Shi et al., 2025), robotic manipulation (Lu et al., 2025b), and embodied navigation (Kim et al., 2025). The rapid advancement of diffusion-based generative models has further spurred this RL-driven innovation wave. Works like ImageReward (Xu et al., 2023) have introduced reinforcement learning to enhance the semantic alignment and visual quality of generative outputs, refining diffusion-based generation through iterative feedback mechanisms derived from human preferences or automated reward critics. Moreover, unified models that blend multiple tasks, i.e., understanding, and generation, into single architectures (Mao et al., 2025; Jiang et al., 2025b) have increasingly relied on RL-driven fine-tuning, achieving generalization and task transfer previously considered challenging. Despite the substantial progress in integrating reinforcement learning with multimodal large language models, several core challenges remain. These include stabilizing policy optimization under complex reward signals, managing high-dimensional and diverse visual inputs, and designing scalable reward functions that support long-horizon decision-making. Addressing these challenges necessitates methodological innovations in both algorithmic design and evaluation protocols.
@@ 3 Reinforcement Learning in Vision
Frameworks(§3.1.1), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  RePIC (Oh et al., 2025), GoalLadder (Zakharov & Whiteson, 2025), Drive-R1 (Li et al., 2025e),
VLM-R1 (Shen et al., 2025b), GRPO-CARE (Chen et al., 2025b), Q-Ponder (Cai et al., 2025c),
@@
ProxyThinker (Xiao et al., 2025b), Jigsaw-R1 (Wang et al., 2025k), SRPO (Wan et al., 2025),
R1-Onevision (Yang et al., 2025f), Reason-RFT (Tan et al., 2025a), Segagent (Zhu et al., 2025c),
VisualQuality-R1 (Wu et al., 2025c), Zhang et al. (Zhang et al., 2025a), SATORI-R1 (Shen et al., 2025a),
Skywork r1v2 (Wang et al., 2025e), Yang et al. (Yang et al., 2025a), etc. , cause_leaf, text width=38em ] ] [ Spatial & 3D Perception
@@
Bindgpt (Zholus et al., 2025), Scene-R1 (Yuan et al., 2025c), Perception-R1 (Yu et al., 2025),
VisRL (Chen et al., 2025h), ViCrit (Wang et al., 2025g), VisionReasoner (Liu et al., 2025e),
ViGoRL (Sarch et al., 2025), Visual-RFT (Liu et al., 2025f), etc. , cause_leaf, text width=38em ] ] [ Image Reasoning (§3.1.3), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Think about Image, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g.,  SVQA-R1 (Wang & Ling, 2025), VL-GenRM (Zhang et al., 2025d), RACRO (Gou et al., 2025)
, EasyARC (Unsal & Akkus, 2025), STAR-R1 (Li et al., 2025g), Visionary-R1 (Xia et al., 2025),
UniVG-R1 (Bai et al., 2025), EchoInk-R1 (Xing et al., 2025), WeThink (Yang et al., 2025b),
@@
Mm-eureka (Meng et al., 2025), Perception-R1 (Xiao et al., 2025a), MiMo-VL (Xiaomi, 2025),
SearchExpert (Li et al., 2025b), etc. , cause_leaf, text width=38em ] ] [ Think with Image, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g.,  Visual Planning (Xu et al., 2025b), GRIT (Fan et al., 2025), VILASR (Wu et al., 2025a),
@@
DeepEyes (Zheng et al., 2025), TACO (Kan et al., 2025), VRAG-RL (Wang et al., 2025f),
VisTA (Huang et al., 2025),TGI (Chern et al., 2025), Chain-of-Focus (Zhang et al., 2025g),
Openthinkimg (Su et al., 2025b), Active-O3 (Zhu et al., 2025d), etc. , cause_leaf, text width=38em ] ] ] [ Video Reasoning(§3.1.4), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  VQ-Insight (Zhang et al., 2025h), TW-GRPO (Dang et al., 2025), EgoVLM (Vinod et al., 2025),
VAU-R1 (Zhu et al., 2025b), DeepVideo-R (Park et al., 2025a), TimeMaster (Zhang et al., 2025e),
VideoR1 (Feng et al., 2025), VideoChat-R1 (Li et al., 2025d), Temporal-RLT (Li et al., 2025a), etc. , cause_leaf, text width=38em ] ] ] [ Visual Generation(§3.2), color=lightplum!100, fill=lightplum!100, text=black [ Image Generation(§3.2.1), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  ImageReward (Xu et al., 2023), ReasonGen-R1 (Zhang et al., 2025i), FocusDiff (Pan et al., 2025),
@@
Black et al., (Black et al., 2023), DiffusionDPO (Wallace et al., 2024), Diffusion-KTO (Li et al., 2024b),
DiffusionRPO (Gu et al., 2024), Miao et al., (Miao et al., 2024b), Parrot (Lee et al., 2024),
DPG-T2I (Wei et al., 2024), RPO (Miao et al., 2024a), RePrompt (Wu et al., 2025b),
RLCM (Oertell et al., 2024), B2-DiffuRL (Hu et al., 2025a), DiffPPO (Xiao et al., 2024),
Simplear(Wang et al., 2025c), LLaVA-Reward (Zhou et al., 2025b), Flow-grpo (Liu et al., 2025a),
Gallici et al. (Gallici & Borde, 2025), TexForce (Chen et al., 2024a), etc. , detect_leaf, text width=38em ] ] [ Video Generation(§3.2.2), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  DanceGRPO (Xue et al., 2025), InfLVG (Fang et al., 2025), Phys-AR (Lin et al., 2025),
VideoReward (Liu et al., 2025b), TeViR (Chen et al., 2025d), GAPO (Zhu et al., 2025a),
Instructvideo (Yuan et al., 2024), Videoscore (He et al., 2024b),
@@
Nabla-R2D3 (Liu et al., 2025d), etc. , detect_leaf, text width=38em ] ] ] [ Unified Model(§3.3), color=harvestgold!100, fill=harvestgold!100, text=black [ Unified RL(§3.3.1), color=harvestgold!100, fill=harvestgold!60, text=black [ e.g.,  UniRL (Mao et al., 2025), CoRL (Jiang et al., 2025b), SelfTok (Wang et al., 2025a),
Hermesflow (Yang et al., 2025c), UnifiedReward (Wang et al., 2025j),
UnifiedReward-Think (Wang et al., 2025i), etc. , detect_harvestgold, text width=38em ] ] [ Task-specific RL(§3.3.2), color=harvestgold!100, fill=harvestgold!60, text=black [ e.g.,  Vargpt-v1. 1 (Zhuang et al., 2025), Emu3 (Wang et al., 2024b), X-Omni (Geng et al., 2025), etc. , detect_harvestgold, text width=38em ] ] ] [ Vision Language Action
@@
UIShift (Gao et al., 2025b), AgentCPM-GUI (Zhang et al., 2025j), MobileGUI-RL (Shi et al., 2025),
ProgRM (Zhang et al., 2025b), Mobile-R1 (Gu et al., 2025), GTA1 (Yang et al., 2025e),
@@
VLN-R1 (Qi et al., 2025), Flare (Hu et al., 2024a), etc. , detect_burlywood, text width=38em ] ] [ Visual Manipulation(§3.4.3), color=burlywood!100, fill=burlywood!60, text=black [ e.g.,  TGRPO (Chen et al., 2025g), RFTF (Shu et al., 2025), RLVLA (Liu et al., 2025c),
@@
RIPT-VLA (Tan et al., 2025b), ReinBot (Zhang et al., 2025c),
@@ 3.1.1 Conventional RL-based MLLMs
We refer to conventional RL-based MLLMs as approaches that apply reinforcement learning primarily to align a vision–language backbone with verifiable, task-level rewards, without explicitly modeling multi-step chain-of-thought reasoning. Typical works RePIC (Oh et al., 2025), GoalLadder (Zakharov & Whiteson, 2025), Drive-R1 (Li et al., 2025e) and VLM-R1 (Shen et al., 2025b) replace preference models with deterministic validators (e.g., exact-match, IoU, BLEU) and optimize the policy by GRPO/PPO variants under a KL regulariser. This design yields stable value-free training, improves zero-shot robustness on captioning, grounding and autonomous-driving benchmarks, and substantially reduces the annotation cost typically incurred by supervised fine-tuning.
@@ 3.1.3 Image Reasoning
Think about Image. Early think about image works for spatial VQA employs view-consistent or transformation-invariant objectives, such as SVQA-R1 (Wang & Ling, 2025) and STAR-R1 (Li et al., 2025g). VL-GenRM (Zhang et al., 2025d) and RACRO (Gou et al., 2025) refine preference data or caption rewards to curb hallucinations. Benchmark-oriented efforts such as EasyARC (Unsal & Akkus, 2025) offer procedurally generated, fully verifiable tasks that suit outcome-based reinforcement learning. To mitigate shortcut reliance and improve generality, Visionary-R1 (Xia et al., 2025) enforces image interpretation before reasoning, whereas UniVG-R1 (Bai et al., 2025) unifies referring, captioning, and detection by coupling a grounding corpus with GRPO fine-tuning. Extensions such as EchoInk-R1 (Xing et al., 2025) further enrich visual reasoning by integrating audio–visual synchrony under GRPO optimization. Meanwhile, curriculum-driven frameworks—WeThink, G1, GThinker, and Observe-R1 progressively increase task complexity or introduce re-thinking cues (e.g., difficulty ladders, multimodal format constraints) to cultivate deeper and more structured reasoning capabilities in MLLMs. These methods show that language only RL with well-designed visual correctness rewards significantly improves model accuracy, robustness, and out-of-distribution performance.
@@ 3.1.4 Video Reasoning
Video reasoning extends the capabilities of MLLMs to process temporal dynamics, requiring not only spatial perception but also sequential understanding and causal inference. Recent works in this domain have proposed diverse approaches to tackle complex reasoning over video inputs. For instance, VQ-Insight (Zhang et al., 2025h) introduces a hierarchical reward design and self-consistency voting mechanism tailored to the question–answering process over long videos. TW-GRPO (Dang et al., 2025) combines token wise credit assignment with GRPO-style optimization to improve fine-grained temporal alignment between textual reasoning and video evidence. Meanwhile, several R1-style frameworks have been developed to unlock video understanding in complex real-world or egocentric settings. EgoVLM (Vinod et al., 2025) and VAU-R1 (Zhu et al., 2025b) focus on egocentric video reasoning with visual memory and utility-based rewards. DeepVideo-R (Park et al., 2025a) integrates dense video encodings and external reward functions to supervise long-horizon reasoning. TimeMaster (Zhang et al., 2025e) explicitly structures temporal abstraction and reasoning via curriculum learning, while VideoR1 (Feng et al., 2025) proposes a scalable RL framework for video-based QA tasks across multiple domains. Collectively, these works highlight the importance of aligning temporal representations with language trajectories through reinforcement learning, paving the way for robust and generalizable video reasoning agents.
@@ 3.2 Visual Generation
Figure 5: Three reward paradigms for RL-based image generation. (a) Human-Centric Preference Optimization: aligns outputs with human aesthetic scores (HPS (Wu et al., 2023b), ImageReward (Xu et al., 2023)); (b) Multimodal Reasoning-Based Evaluation: scores images via multimodal reasoning consistency (UnifiedReward (Wang et al., 2025j), PARM (Guo et al., 2025d)); (c) Metric-Driven Objective Optimization: minimizes task-specific quantitative metrics such as FID and IoU.